Which of the following equations define a plane? 2 ド!..lil..lil,r..e R X-H-N+3하a,JER
Answer #1

All planes are of the form X=a+t.b+s.c where t,s€|R.Then the panle passes through a and parallel to both b and c.Hence X represents a plane if b and c are parallel.For option (b) we see that -1/2=2/(-4)=1/(-2).Therefore [-1 2 1]^T and [2 -4 -2]^T are parallel.Therefore option (b) represents a plane.